# Overcomplete Tensor Decomposition via Koszul–Young Flattenings

### v1

@@ 2.1.2 Genericity and Symbolic Rank
Throughout, we use the standard meaning of the term “generic” from algebraic geometry. Namely, for symbolic variables $x=(x_{1},\ldots,x_{p})$, we say that a property holds true “generically” (or holds for “generically chosen $x$”) if there exists a nonidentically-zero polynomial $P(x_{1},\ldots,x_{p})\in\mathbb{R}[x]$ such that the property holds true for all $x\in\mathbb{R}^{n}$ such that $P(x)\neq 0$. If a property holds true generically, it holds for “almost all” $x$ in the sense that the set of $x$ for which the property fails is measure zero. If a finite collection of properties each holds true generically, then it is also generically true that all the properties hold simultaneously; this can be seen by multiplying together all the associated polynomials $P(x)$.
@@ 2.2.2 Koszul–Young Flattenings
Our new algorithm is inspired by an idea first introduced by [LO13] as “Young flattenings” [LO13, Lan11, Lan15] and later referred to as “Koszul flattenings” [OO13, LM17, HOOS19] or “Koszul–Young flattenings” [Far16, Gua17]. Essentially, Koszul–Young flattenings provide a non-trivial linear map from tensors to matrices. These ideas have already appeared in tensor decomposition algorithms [OO13], but for $n\times n\times n$ tensors, these results only handle the undercomplete case $r\leq n$. More recently, [Per18] used Koszul–Young flattenings to solve rank detection for overcomplete tensors with generic components, namely $n\times n\times n$ tensors of rank up to $23n/$ 2. Furthermore, [Per18] proposed a hierarchy of poly-time approaches for rank detection that conjecturally approaches rank $2n$, as well as an approach for decomposition that conjecturally tolerates rank $(1+\epsilon)n$ for a small constant $\epsilon>0$. We will resolve and extend these conjectures: for sufficiently large $n\times n\times n$ tensors with generic components, we solve both rank detection and decomposition in polynomial time, up to rank $(2-\epsilon)n$ for an arbitrary constant $\epsilon>0$. We will also handle non-square tensors.
While the Koszul–Young flattenings are typically described in abstract algebraic terms, we will give instead a very explicit combinatorial description of a matrix based on these ideas. Suppose $T$ is $n_{1}\times n_{2}\times n_{3}$ and fix integers $p,q$ with $0\leq p\leq p+1\leq q\leq n_{1}$. Define a matrix $M=M(T;p,q)$ with rows indexed by
@@
The dimensions of $M$ are $\binom{q}{p}n_{2}\times\binom{q}{p+1}n_{3}$. We will leave $p,q$ as free variables for now, but eventually it will be advantageous to take $q$ to be a (large) constant so that $M$ has polynomial size, and $p\approx q\frac{n_{3}}{n_{2}+n_{3}}$ which makes $M$ as square as possible. One special case of interest is $n_{2}=n_{3}$, in which case we can take $q$ odd with $q=2p+1$ so that $M$ is exactly square, since $\binom{2p+1}{p}=\binom{2p+1}{p+1}$; the prior work [Per18] considers only this square case and mainly focuses on $p=1$. We remark that the extreme cases $p=0$ and $p=q-1$ essentially reduce to “trivial” flattenings from the previous section (but with additional signs in the case $p=q-1$).
@@ 2.2.3 Rank Detection
Therefore, as long as neither $b$ nor $c$ is the zero vector, $M$ has the same rank as $A$. We will see that the matrix $A$ is rank-deficient, with generic rank $\binom{q-1}{p}$. If, for instance, $q=2p+1$ then $A$ is square since $\binom{2p+1}{p}=\binom{2p+1}{p+1}$, but crucially, its generic rank $\binom{2p}{p}$ is roughly half its dimension (for large $p$). The following is proved in Section 3.1.
Proposition 2.3.
@@
For an arbitrary tensor $T$, we always have $\mathrm{rank}(M(T;p,q))\leq\binom{q-1}{p}\cdot\mathrm{rank}(T)$. This follows from the fact that the rank of $A$ from (6) never exceeds its generic rank $\binom{q-1}{p}$; see Lemma 3.1. As a result, $M$ provides a certificate that $\mathrm{rank}(T)\geq\mathrm{rank}(M(T;p,q))/\binom{q-1}{p}$ for any given tensor $T$.
@@ 2.2.4 Decomposition Algorithm
To recap, we can compute the left-hand side of (10) from the flattening $M$, which gives us access to the right-hand side. This means we know the span of $r$ different rank-1 tensors $d^{(\ell)}\otimes b^{(\ell)}$ (which can also be viewed as rank-1 matrices of dimension $(p+1)\times n_{2}$) and our next objective will be to extract the individual rank-1 elements $d^{(\ell)}\otimes b^{(\ell)}$. Since our rank-1 tensors are generic, the work of [JLV23] provides an algorithm for precisely this task, under the condition $r\leq p(n_{2}-1)/4$; see Corollary 3 of [JLV23], which considers a generalization of this task (in their notation, we only need the case $s=R$). This allows us to recover vectors $\hat{d}^{(\ell)}$ and $\hat{b}^{(\ell)}$ for $\ell\in[r]$, such that $\hat{d}^{(\ell)}=\alpha_{\ell}\cdot d^{(\pi(\ell))}$ and $\hat{b}^{(\ell)}=\beta_{\ell}\cdot b^{(\pi(\ell))}$ where $\alpha_{\ell},\beta_{\ell}\in\mathbb{R}$ are nonzero scalars and $\pi$ is a permutation $[r]\to[r]$.
At this point, we repeat the entire algorithm so far with the “$b$” and “$c$” modes switched. We use the same value for $q$, but use $q-p-1$ in place of $p$ so that the new flattening remains as square as possible. This allows us to obtain vectors $\hat{c}^{(\ell)}$ and $\hat{f}^{(\ell)}$ which serve as estimates for $c^{(\ell)}$ and
$f^{(\ell)}:=(a^{(\ell)}_{1},\ldots,a^{(\ell)}_{q-p})^{\top},$
@@
The full algorithm is presented as Algorithm 1 below. The second phase of the algorithm involves a flattening $M^{\prime}=M^{\prime}(T;p,q)$ that is identical to $M$ but with the “$b$” and “$c$” modes switched, and with $q-p-1$ in place of $p$. Formally, $M^{\prime}(T;p,q)$ is indexed by
@@
We have assumed $\bar{p}\geq 2$, since otherwise Step 7 will not have a unique solution. If $\bar{p}=1$, this means either $p=0$ or $p=q-1$, which are the extreme cases where $M(T;p,q)$ essentially reduces to a trivial flattening. In these cases, one should use a simpler decomposition algorithm of [JLV23] based on the trivial flattening, as we mentioned in Section 2.2.1; this algorithm works under the condition (3).
@@ 2.2.5 Uniqueness Theorem and Algorithmic Guarantees
Some of the conditions below may appear rather opaque, so we first briefly motivate what role they play. It is natural for the matrix $M$ to appear (condition (vi)), given its use as a rank detection device (Theorem 2.4) and its role in the decomposition algorithm, but we will also define two additional matrices $N$ and $P$. The condition (viii) on $N$ is used to ensure that (8) holds (no spurious elements in the subspace intersection), while the condition (x) on $P$ is used to ensure that the [JLV23] algorithm succeeds. The matrices $M^{\prime},N^{\prime},P^{\prime}$ are identical to $M,N,P$ but with the “$b$” and “$c$” modes of the tensor switched and with $p$ replaced by $q-p-1$.
@@
    The $r$ vectors $(a^{(\ell)}_{1},\ldots,a^{(\ell)}_{\bar{p}})^{\top}$ for $\ell\in[r]$ are pairwise linearly independent, i.e., no one of these vectors is a scalar multiple of another.
@@
    The $r$ vectors $\{d^{(\ell)}\otimes b^{(\ell)}\,:\,\ell\in[r]\}$ are linearly independent, where $d^{(\ell)}:=(a^{(\ell)}_{1},\ldots,a^{(\ell)}_{p+1})^{\top}$.
@@
    The $r$ vectors $\{f^{(\ell)}\otimes c^{(\ell)}\,:\,\ell\in[r]\}$ are linearly independent, where $f^{(\ell)}:=(a^{(\ell)}_{1},\ldots,a^{(\ell)}_{q-p})^{\top}$.
@@
The proof can be found in Section 3.6. This is qualitatively the same behavior as our rank detection result. Suppose $n_{1}\leq n_{2}\leq n_{3}$ and consider the regime $n_{1}\to\infty$ with $\alpha$ of constant order. For any constant $\epsilon>0$ we can handle $r$ as large as $(1-\epsilon)(n_{2}+n_{3})$ by taking $q$ to be a sufficiently large constant (depending on $\alpha,\epsilon$). As long as $q$ is a constant, the runtime of Algorithm 1 is polynomial in the largest dimension $n_{3}$.
@@ 2.3.1 Flattenings of Koszul–Young Type
The proof can be found in Section 4.2. Recall from Section 2.1.2 that symbolic matrices have a well-defined notion of generic rank, so the expressions on the left- and right-hand sides of (19) must each take one specific value when the components are chosen generically. (We cannot have a situation where some positive-measure set of components leads to rank $r_{1}$ while some other positive-measure set of components leads to a different rank $r_{2}$.) The right-hand side of (19) could equivalently be written as
@@ 2.3.2 Linear Flattenings
The proof can be found in Section 4.3. This result is similar to Theorem 4.4 of [EGOW18], where the bound $r\leq 8n$ is proven for “square” tensors, i.e., tensors of format $n\times n\times n$. For square tensors, our result gives an improvement from $8n$ to $26n+$ 2. Our argument is similar to that of [EGOW18] and we describe the difference in Remark 4.2.
@@ 2.4 Open Problems
  • For an $n\times n\times n$ tensor with generic components, what is the maximum rank for which a linear flattening can have additivity? We have shown the answer is asymptotically $cn$ for a constant $c\in[2,6]$. This improved the previous range of $[3/2,8]$ by [Per18] and [EGOW18], but the exact value of $c$ remains unknown.
@@ 3 Proofs for Upper Bounds
Fix $i\in[q]$ such that $a_{i}\neq 0$. We first show $\mathrm{rank}(A)\geq\binom{q-1}{p}$ by demonstrating a $\binom{q-1}{p}\times\binom{q-1}{p}$ submatrix with full rank. Specifically, consider the submatrix indexed by $\{S\,:\,i\notin S\}\times\{U\,:\,i\in U\}$. Note that this submatrix has exactly one nonzero entry per row and exactly one nonzero entry per column, and all the nonzero entries are equal to ±$\pm a_{i}$. This shows $\mathrm{rank}(A)\geq\binom{q-1}{p}$.
@@
Provided $a_{i}\neq 0$ for all $i\in[q]$ (which is true for generic $a$), the matrix $A=A(a;p,q)$ defined in (6) can be factored as $A=\mathrm{diag}(v)\cdot\tilde{A}\cdot\mathrm{diag}(w)$ where $\tilde{A}_{SU}=\sum_{i\in[q]}\mathbbm{1}_{U=S\sqcup\{i\}}\cdot\sigma(U,i)$, $v_{S}=\prod_{i\in S}a_{i}^{-1}$, and $w_{U}=\prod_{i\in U}a_{i}$. Note that $\tilde{A}=A(\mathbbm{1})$ where $\mathbbm{1}:=(1,\ldots,1)^{\top}$, so by Lemma 3.1 we have $\mathrm{rank}(\tilde{A})=\binom{q-1}{p}$. We can therefore factor $\tilde{A}=\tilde{Q}\tilde{R}^{\top}$ where $\tilde{Q},\tilde{R}$ are $\binom{q}{p}\times\binom{q-1}{p}$ and $\binom{q}{p+1}\times\binom{q-1}{p}$, respectively. Now write $A=QR^{\top}$ where $Q=\mathrm{diag}(v)\cdot\tilde{Q}$ and $R=\mathrm{diag}(w)\cdot\tilde{R}$. We use the superscript $(\ell)$ to mark that a matrix is built from the tensor component $a^{(\ell)}$, so for instance, $A^{(\ell)}:=A(a^{(\ell)};p,q)$. Factor $M$ as
@@
where, e.g., $P^{(1)}\otimes b^{(1)}$ denotes the $\binom{q}{p}n_{2}\times\binom{q-1}{p}$ Kronecker product. This gives a factorization of $M$ as the product of two matrices with inner dimension $r\binom{q-1}{p}$, so it suffices to show that the two factors on the right-hand side of (20) have full column rank and full row rank, respectively. We will show this for the first factor and then explain the (minor) changes needed to adapt the proof for the second factor.
@@
We aim to show $\det(Q^{\prime\prime})$ is nonzero as a polynomial in the variables $b_{j\ell},a_{i\ell}^{-1}$. In the usual expansion of $\det(P^{\prime\prime})$ as a sum over permutations, each term chooses exactly 1 entry per row and column of $Q^{\prime\prime}$. This choice can be broken down as first choosing a set of cells $(S,j,\ell)$ with exactly 1 cell per row $(S,j)$ and exactly $\binom{q-1}{p}$ cells per “pillar” $\ell$ (where “pillar” describes a vertical stack of cells that are all indexed by the same $\ell$ value), and then choosing one entry per cell in a way that each column appears once. Let $H$ denote a subset of tuples $(S,j,\ell)$ representing the choice at the first stage. For a given $H$, define matrices $F_{1}^{H},\ldots,F_{q}^{H}$, each of size $\binom{q-1}{p}\times\binom{q-1}{p}$, where $F_{\ell}^{H}$ is formed by collecting the rows $\tilde{Q}_{S}$ for each $(S,j,\ell)\in H$ (i.e., the cells selected in pillar $\ell$). Now
@@
To complete the proof, we will identify a particular monomial that has nonzero coefficient in $\det(Q^{\prime\prime})$. To this end, define a “good” set of cells $G$ as follows, which (like $H$) will have 1 cell per row and $\binom{q-1}{p}$ cells per pillar. Within each pillar $\ell\in[q]$, for each $S$ with $\ell\notin S$, include the cell $(S,j,\ell)$ where $j$ is the first available index (i.e., the minimum $j\in[q-p]$ such that $(S,j,\ell^{\prime})$ has not already been declared a good cell for some $\ell^{\prime}<\ell$). Each $S$ is used $q-p$ times in total (across all pillars), so the $j$ values will not “run out.” We claim that the only term in (21) that produces the monomial $(a,b)^{G}$ (with nonzero coefficient) is $H=G$: within each pillar $\ell=1,2,\ldots$, the powers of $a$ (particularly the absence of $a_{\ell\ell}^{-1}$) determine the collection of $S$ values that $H$ must use (namely one copy of each $S$ for which $\ell\notin S$), and the powers of $b$ determine the multiset of $j$ values to use. The choice of “first available $j$” used in $G$ constrains a unique pairing between $S$ and $j$ values.
We conclude that the coefficient of $(a,b)^{G}$ in $\det(Q^{\prime\prime})$ is ±$\pm\prod_{\ell=1}^{q}\det(F_{\ell}^{G})$ and it remains to show that each $\det(F_{\ell}^{G})$ is nonzero. This follows because $F_{\ell}^{G}$ has rows $\tilde{Q}_{S}$ for all $S$ with $\ell\notin S$, and the corresponding rows of $\tilde{A}=A(\mathbbm{1})=\tilde{Q}\tilde{R}^{\top}$ are linearly independent by Lemma 3.1.
@@
Define $\tilde{d}^{(\ell)}:=(\tilde{a}^{(\ell)}_{1},\ldots,\tilde{a}^{(\ell)}_{p+1})^{\top}$. From (26) we have $z^{(V)}(\tilde{a}^{(\ell)};p,q)\otimes\tilde{b}^{(\ell)}\in\mathrm{colspan}(M)\cap(Z_{p,q}\otimes\mathbb{R}^{n_{2}})$ for all $\ell\in[r]$, and so $\tilde{d}^{(\ell)}\otimes\tilde{b}^{(\ell)}\in(\varphi_{p,q}\otimes I)(\mathrm{colspan}(M)\cap(Z_{p,q}\otimes\mathbb{R}^{n_{2}}))$. But by (27) and Lemma 3.3, the only rank-1 tensors in this subspace are scalar multiples of $d^{(\ell)}\otimes b^{(\ell)}$. We therefore conclude that for every $\ell\in[r]$ there exists $\ell^{\prime}\in[r]$ such that $\tilde{d}^{(\ell)}\otimes\tilde{b}^{(\ell)}$ is a scalar multiple of $d^{(\ell^{\prime})}\otimes b^{(\ell^{\prime})}$.
Next we need to repeat the argument with the “$b$” and “$c$” modes switched and with $p$ replaced by $q-p-1$, which will use the assumptions (iv), (vii), (ix), (xi). Due to symmetry, the argument is identical to before, and we only state the main claims. First, the flattening $M^{\prime}$ takes the form
$M^{\prime}(T;p,q)=\sum_{\ell=1}^{r}A(a^{(\ell)};q-p-1,q)\otimes(c^{(\ell)}b^{(\ell)\top})=\sum_{\ell=1}^{r}A(\tilde{a}^{(\ell)};q-p-1,q)\otimes(\tilde{c}^{(\ell)}\tilde{b}^{(\ell)\top}),$
and by Lemma 3.1, the maximum possible rank of $A(a;q-p-1,q)$ is $\binom{q-1}{q-p-1}=\binom{q-1}{p}$. Analogous to (25) and (26), we have
@@
where, recall, $f^{(\ell)}:=(a^{(\ell)}_{1},\ldots,a^{(\ell)}_{q-p})^{\top}$. Analogous to Lemma 3.3, the vectors $\{f^{(\ell)}\otimes c^{(\ell)}\,:\,\ell\in[r]\}$ are the only rank-1 tensors in their span, up to scalar multiple, and the algorithm of [JLV23] recovers them from their span. Finally, defining $\tilde{f}^{(\ell)}:=(\tilde{a}^{(\ell)}_{1},\ldots,\tilde{a}^{(\ell)}_{q-p})^{\top}$, we have $\tilde{f}^{(\ell)}\otimes\tilde{c}^{(\ell)}\in(\varphi_{q-p-1,q}\otimes I)(\mathrm{colspan}(M^{\prime})\cap(Z_{q-p-1,q}\otimes\mathbb{R}^{n_{3}}))$ and so every $\tilde{f}^{(\ell)}\otimes\tilde{c}^{(\ell)}$ is a scalar multiple of some $f^{(\ell^{\prime\prime})}\otimes c^{(\ell^{\prime\prime})}$.

### v2

@@ 2.1.2 Genericity and Symbolic Rank
Throughout, we use the standard meaning of the term “generic” from algebraic geometry. Namely, for symbolic variables $x=(x_{1},\ldots,x_{p})$, we say that a property holds true “generically” (or holds for “generically chosen xx”) if there exists a nonidentically-zero polynomial $P(x_{1},\ldots,x_{p})\in\mathbb{R}[x]$ such that the property holds true for all $x\in\mathbb{R}^{n}$ such that $P(x)\neq 0$. If a property holds true generically, it holds for “almost all” $x$ in the sense that the set of $x$ for which the property fails is measure zero. If a finite collection of properties each holds true generically, then it is also generically true that all the properties hold simultaneously; this can be seen by multiplying together all the associated polynomials $P(x)$.
@@ 2.2.2 Koszul–Young Flattenings
Our new algorithm is inspired by an idea first introduced by [LO13] as “Young flattenings” [LO13, Lan11, Lan15] and later referred to as “Koszul flattenings” [OO13, LM17, HOOS19] or “Koszul–Young flattenings” [Far16, Gua17]. (We thank J.M. Landsberg for clarifying that Koszul flattenings are a special case of Young flattenings, with the optimal choice of parameters.) Essentially, Koszul–Young flattenings provide a non-trivial linear map from tensors to matrices. These ideas have already appeared in tensor decomposition algorithms [OO13], but for $n\times n\times n$ tensors, these results only handle the undercomplete case $r\leq n$. More recently, [Per18] used Koszul–Young flattenings to solve rank detection for overcomplete tensors with generic components, namely $n\times n\times n$ tensors of rank up to $3n/2$. Furthermore, [Per18] proposed a hierarchy of poly-time approaches for rank detection that conjecturally approaches rank $2n$, as well as an approach for decomposition that conjecturally tolerates rank $(1+\epsilon)n$ for a small constant $\epsilon>0$. We will resolve and extend these conjectures: for sufficiently large $n\times n\times n$ tensors with generic components, we solve both rank detection and decomposition in polynomial time, up to rank $(2-\epsilon)n$ for an arbitrary constant $\epsilon>0$. We will also handle non-square tensors.
The Koszul–Young flattenings are typically described in abstract algebraic terms, namely as a map $B^{*}\otimes\wedge^{p}(A)\rightarrow\wedge^{p+1}A\otimes C$. We will give instead a very explicit combinatorial description of a matrix based on these ideas. Our construction has two parameters $p,q$ instead of only $p$, and specializes to (the transpose of) the usual Koszul–Young flattenings when $q=2p+1$. Suppose $T$ is $n_{1}\times n_{2}\times n_{3}$ and fix integers $p,q$ with $0\leq p<p+1\leq q\leq n_{1}$. Define a matrix $M=M(T;p,q)$ with rows indexed by
@@
The dimensions of $M$ are $\binom{q}{p}n_{2}\times\binom{q}{p+1}n_{3}$. We will leave $p,q$ as free variables for now, but eventually it will be advantageous to take $q$ to be a (large) constant so that $M$ has polynomial size, and $p\approx q\frac{n_{3}}{n_{2}+n_{3}}$ which makes $M$ as square as possible. One special case of interest is $n_{2}=n_{3}$, in which case we can take $q$ odd with $q=2p+1$ so that $M$ is exactly square, since $\binom{2p+1}{p}=\binom{2p+1}{p+1}$; the prior work [Per18] considers only this square case and mainly focuses on $p=1$. We remark that the extreme cases $p=0$ and p=q1p=q-1 essentially reduce to “trivial” flattenings from the previous section (but with additional signs in the case p=q1p=q-1).
@@ 2.2.3 Rank Detection
Therefore, as long as neither $b$ nor $c$ is the zero vector, $M$ has the same rank as $A$. We will see that the matrix $A$ is rank-deficient, with generic rank $\binom{q-1}{p}$. If, for instance, $q=2p+1$ then $A$ is square since $\binom{2p+1}{p}=\binom{2p+1}{p+1}$, but crucially, its generic rank $\binom{2p}{p}$ is roughly half its dimension (for large $p$). This is a known fact, and we provide a proof in Section 3.1.
Proposition 2.3 ([LO13]).
@@
How does the runtime scale with $\epsilon$? Since $q$ is a constant, it suffices to have $(1+\alpha)/q\leq\epsilon/2$, meaning we should choose $q\geq 2(1+\alpha)/\epsilon$. Now each dimension of $M$ is bounded by $n_{3}\cdot 2^{q}=n_{3}\cdot 2^{O(\alpha/\epsilon)}$. The runtime is therefore polynomial in $n_{3}\cdot 2^{\alpha/\epsilon}$. It might be possible to remove the dependence on $\alpha$ with a more careful analysis.
@@
For an arbitrary tensor $T$, we always have $\mathrm{rank}(M(T;p,q))\leq\binom{q-1}{p}\cdot\mathrm{rank}(T)$. This follows from the fact that the rank of $A$ from (6) never exceeds its generic rank $\binom{q-1}{p}$; see Lemma 3.1. As a result, $M$ provides a certificate that $\mathrm{rank}(T)\geq\lceil\mathrm{rank}(M(T;p,q))/\binom{q-1}{p}\rceil$ for any given tensor $T$.
@@ 2.2.4 Decomposition Algorithm
To recap, we can compute the left-hand side of (10) from the flattening $M$, which gives us access to the right-hand side. This means we know the span of $r$ different rank-1 tensors $d^{(\ell)}\otimes b^{(\ell)}$ (which can also be viewed as rank-1 matrices of dimension $(p+1)\times n_{2}$) and our next objective will be to extract the individual rank-1 elements $d^{(\ell)}\otimes b^{(\ell)}$. Since our rank-1 tensors are generic, the work of [JLV23] provides an algorithm for precisely this task, under the condition $r\leq p(n_{2}-1)/4$; see Corollary 3 of [JLV23], which considers a generalization of this task (in their notation, we only need the case $s=R$). This allows us to recover vectors $\hat{d}^{(\ell)}$ and $\hat{b}^{(\ell)}$ for $\ell\in[r]$, such that $\hat{d}^{(\ell)}=\alpha_{\ell}\cdot d^{(\pi(\ell))}$ and $\hat{b}^{(\ell)}=\beta_{\ell}\cdot b^{(\pi(\ell))}$ where $\alpha_{\ell},\beta_{\ell}\in\mathbb{R}$ are nonzero scalars and $\pi$ is a permutation $r]\to[r]$.
At this point, we repeat the entire algorithm so far with the “bb” and “cc” modes switched. We use the same value for $q$, but use qp1q-p-1 in place of $p$ so that the new flattening remains as square as possible. This allows us to obtain vectors $\hat{c}^{(\ell)}$ and $\hat{f}^{(\ell)}$ which serve as estimates for $c^{(\ell)}$ and
[…]
@@
The full algorithm is presented as Algorithm 1 below. The second phase of the algorithm involves a flattening $M^{\prime}=M^{\prime}(T;p,q)$ that is identical to $M$ but with the “bb” and “cc” modes switched, and with qp1q-p-1 in place of $p$. Formally, $M^{\prime}(T;p,q)$ is indexed by
@@
We have assumed $\bar{p}\geq 2$, since otherwise Step 7 will not have a unique solution. If $\bar{p}=1$, this means either $p=0$ or p=q1p=q-1, which are the extreme cases where $M(T;p,q)$ essentially reduces to a trivial flattening. In these cases, one should use a simpler decomposition algorithm of [JLV23] based on the trivial flattening, as we mentioned in Section 2.2.1; this algorithm works under the condition (3).
@@ 2.2.5 Uniqueness Theorem and Algorithmic Guarantees
Some of the conditions below may appear rather opaque, so we first briefly motivate what role they play. It is natural for the matrix $M$ to appear (condition (vi)), given its use as a rank detection device (Theorem 2.4) and its role in the decomposition algorithm, but we will also define two additional matrices $N$ and $P$. The condition (viii) on $N$ is used to ensure that (8) holds (no spurious elements in the subspace intersection), while the condition (x) on $P$ is used to ensure that the [JLV23] algorithm succeeds. The matrices $M^{\prime},N^{\prime},P^{\prime}$ are identical to $M,N,P$ but with the “bb” and “cc” modes of the tensor switched and with $p$ replaced by qp1q-p-1.
@@
[…]
@@
[…]
@@
[…]
@@
The proof can be found in Section 3.6. This is qualitatively the same behavior as our rank detection result. Suppose $n_{1}\leq n_{2}\leq n_{3}$ and consider the regime $n_{1}\to\infty$ with $\alpha$ of constant order. For any constant $\epsilon>0$ we can handle $r$ as large as $(1-\epsilon)(n_{2}+n_{3})$ by taking $q$ to be a sufficiently large constant (depending on $\alpha,\epsilon$). As long as $q$ is a constant, the runtime of Algorithm 1 is polynomial in the largest dimension $n_{3}$ (more precisely, polynomial in $n_{3}\cdot 2^{\alpha/\epsilon}$).
@@ 2.3.1 Flattenings of Koszul–Young Type
The proof can be found in Section 4.2. We thank an anonymous reviewer for pointing out that the case $n_{2}=n_{3}$ of this result appears in a remark after Theorem 6.1 of [DM18]. Recall from Section 2.1.2 that symbolic matrices have a well-defined notion of generic rank, so the expressions on the left- and right-hand sides of (19) must each take one specific value when the components are chosen generically. (We cannot have a situation where some positive-measure set of components leads to rank $r_{1}$ while some other positive-measure set of components leads to a different rank $r_{2}$.) The right-hand side of (19) could equivalently be written as
@@ 2.3.2 Linear Flattenings
The proof can be found in Section 4.3. This result is similar to Theorem 4.4 of [EGOW18], where the bound $r\leq 8n$ is proven for “square” tensors, i.e., tensors of format $n\times n\times n$. For square tensors, our result gives an improvement from $8n$ to $6n+2$. Our argument is similar to that of [EGOW18] and we describe the difference in Remark 4.2. We thank an anonymous reviewer for pointing out that the improvement to $6n+2$ was previously observed by Visu Makam in the talk [Mak22].
@@ 2.4 Open Problems
  • For an $n\times n\times n$ tensor with generic components, what is the maximum rank for which a linear flattening can have additivity? We have shown the answer is asymptotically cncn for a constant $c\in[2,6]$. This improved the previous range of $[3/2,8]$ by [Per18] and [EGOW18], but the exact value of $c$ remains unknown.
@@ 3 Proofs for Upper Bounds
Fix $i\in[q]$ such that $a_{i}\neq 0$. We first show $\mathrm{rank}(A)\geq\binom{q-1}{p}$ by demonstrating a $\binom{q-1}{p}\times\binom{q-1}{p}$ submatrix with full rank. Specifically, consider the submatrix indexed by $\{S\,:\,i\notin S\}\times\{U\,:\,i\in U\}$. Note that this submatrix has exactly one nonzero entry per row and exactly one nonzero entry per column, and all the nonzero entries are equal to $\pm a_{i}$. This shows $\mathrm{rank}(A)\geq\binom{q-1}{p}$.
@@
Provided $a_{i}\neq 0$ for all $i\in[q]$ (which is true for generic $a$), the matrix $A=A(a;p,q)$ defined in (6) can be factored as $A=\mathrm{diag}(v)\cdot\tilde{A}\cdot\mathrm{diag}(w)$ where $\tilde{A}_{SU}=\sum_{i\in[q]}\mathbbm{1}_{U=S\sqcup\{i\}}\cdot\sigma(U,i)$, $v_{S}=\prod_{i\in S}a_{i}^{-1}$, and $w_{U}=\prod_{i\in U}a_{i}$. Note that $\tilde{A}=A(\mathbbm{1};p,q)$ where $\mathbbm{1}:=(1,\ldots,1)^{\top}$, so by Lemma 3.1 we have $\mathrm{rank}(\tilde{A})=\binom{q-1}{p}$. We can therefore factor $\tilde{A}=\tilde{Q}\tilde{R}^{\top}$ where $\tilde{Q},\tilde{R}$ are $\binom{q}{p}\times\binom{q-1}{p}$ and $\binom{q}{p+1}\times\binom{q-1}{p}$, respectively. Now write $A=QR^{\top}$ where $Q=\mathrm{diag}(v)\cdot\tilde{Q}$ and $R=\mathrm{diag}(w)\cdot\tilde{R}$. We use the superscript $(\ell)$ to mark that a matrix is built from the tensor component $a^{(\ell)}$, so for instance, $A^{(\ell)}:=A(a^{(\ell)};p,q)$. Factor $M$ as
@@
where, e.g., $Q^{(1)}\otimes b^{(1)}$ denotes the $\binom{q}{p}n_{2}\times\binom{q-1}{p}$ Kronecker product. This gives a factorization of $M$ as the product of two matrices with inner dimension $r\binom{q-1}{p}$, so it suffices to show that the two factors on the right-hand side of (20) have full column rank and full row rank, respectively. We will show this for the first factor and then explain the (minor) changes needed to adapt the proof for the second factor.
@@
We aim to show $\det(Q^{\prime\prime})$ is nonzero as a polynomial in the variables $b_{j\ell},a_{i\ell}^{-1}$. In the usual expansion of $\det(Q^{\prime\prime})$ as a sum over permutations, each term chooses exactly 1 entry per row and column of $Q^{\prime\prime}$. This choice can be broken down as first choosing a set of cells $(S,j,\ell)$ with exactly 1 cell per row $(S,j)$ and exactly $\binom{q-1}{p}$ cells per “pillar” $\ell$ (where “pillar” describes a vertical stack of cells that are all indexed by the same $\ell$ value), and then choosing one entry per cell in a way that each column appears once. Let $H$ denote a subset of tuples $(S,j,\ell)$ representing the choice at the first stage. For a given $H$, define matrices $F_{1}^{H},\ldots,F_{q}^{H}$, each of size $\binom{q-1}{p}\times\binom{q-1}{p}$, where $F_{\ell}^{H}$ is formed by collecting the rows $\tilde{Q}_{S}$ for each $(S,j,\ell)\in H$ (i.e., the cells selected in pillar $\ell$). Now
@@
To complete the proof, we will identify a particular monomial that has nonzero coefficient in $\det(Q^{\prime\prime})$. To this end, define a “good” set of cells $G$ as follows, which (like $H$) will have 1 cell per row and $\binom{q-1}{p}$ cells per pillar. Within each pillar $\ell\in[q]$, for each $S$ with $\ell\notin S$, include the cell $(S,j,\ell)$ where $j$ is the first available index (i.e., the minimum $j\in[q-p]$ such that $(S,j,\ell^{\prime})$ has not already been declared a good cell for some $\ell^{\prime}<\ell$). Each $S$ is used qpq-p times in total (across all pillars), so the $j$ values will not “run out.” We claim that the only term in (21) that produces the monomial $(a,b)^{G}$ (with nonzero coefficient) is $H=G$: within each pillar $\ell=1,2,\ldots$, the powers of $a$ (particularly the absence of $a_{\ell\ell}^{-1}$) determine the collection of $S$ values that $H$ must use (namely one copy of each $S$ for which $\ell\notin S$), and the powers of $b$ determine the multiset of $j$ values to use. The choice of “first available jj” used in $G$ constrains a unique pairing between $S$ and $j$ values.
We conclude that the coefficient of $(a,b)^{G}$ in $\det(Q^{\prime\prime})$ is $\pm\prod_{\ell=1}^{q}\det(F_{\ell}^{G})$ and it remains to show that each $\det(F_{\ell}^{G})$ is nonzero. This follows because $F_{\ell}^{G}$ has rows $\tilde{Q}_{S}$ for all $S$ with $\ell\notin S$, and the corresponding rows of $\tilde{A}=A(\mathbbm{1})=\tilde{Q}\tilde{R}^{\top}$ are linearly independent by Lemma 3.1.
@@
[…]
Next we need to repeat the argument with the “bb” and “cc” modes switched and with $p$ replaced by qp1q-p-1, which will use the assumptions (iv), (vii), (ix), (xi). Due to symmetry, the argument is identical to before, and we only state the main claims. First, the flattening $M^{\prime}$ takes the form
[…]
and by Lemma 3.1, the maximum possible rank of A(a;qp1,q)A(a;q-p-1,q) is $\binom{q-1}{q-p-1}=\binom{q-1}{p}$. Analogous to (25) and (26), we have
@@
[…]